OpenStudy (kimberly_pr):

solve for y (4y+4)/3 + (7y-8)/6 = 15

OpenStudy (kg1975):

Substitute 1 in this equation , 4(-4y - 25) - 3y = 33 -16y - 100 - 3y = 33 -19y = 133 y = -133/19 = -7 So , x = 3 from 1

OpenStudy (kimberly_pr):

You mean y+3?

OpenStudy (kg1975):

yea XD sorry

OpenStudy (kg1975):

fast typing XD

OpenStudy (kimberly_pr):

Lol it's ok, thanks

OpenStudy (kg1975):

yw just tag me in any time u need help im open for everyone

OpenStudy (kimberly_pr):

Can you help me with one more? :)

OpenStudy (kg1975):

sure

OpenStudy (kg1975):

u know how to tag ppl right?

OpenStudy (kimberly_pr):

Simplify z^-8/ z^2 And yes I do

OpenStudy (kimberly_pr):

@kg1975

OpenStudy (kimberly_pr):

Would it be z^-10?

OpenStudy (kg1975):

let me see

OpenStudy (kg1975):

i believe so because u add the 8 and the 2 and since the 8 is a - and bigger it becomes and -10 z^-10 you are correct

OpenStudy (kg1975):

@KingNoob

OpenStudy (kimberly_pr):

Yeah because when dividing you subtract, so -8 -2= -10 but I'm not 100% sure

OpenStudy (kg1975):

its right

OpenStudy (kg1975):

im gonna make a room for us so u can post ur stuff its hard finding your problem when i help others

OpenStudy (kg1975):

ok? and ill tag u in the room